Telescope and AGw Commissioning at LFBG with IRTC - 03-04 April 2009 MST

Observers: J. Hill, R. Green (mtn)
Telescope Operator: A. Ceranski
Telescope Support: J. Rousseau

Summary

Winds gusting up to 38 m/sec, light snow in forecast. No observing tonight. We'll do some more software tests.

Details

OSS tests

20:18 Used "Move to FS" button on OSS.M3 GUI to rotate M3 selector. Cool.

20:21 Confirmed Bob's assertion that the selector is accurate to 0.01 degrees.

We still have the slightly strange behaviour that the selector moves and reaches "in-position", and the OSS command completes. However, the selector position continues to move slightly after that for 30-90 sec.

PCS / MCSPU tests

20:30 Aaron goes upstairs to check the power on the RIM module. That was OK. The problem turned out to be the cable between jet and the Reef module that got knocked loose while messing with the GPS.

21:57 pointto BS9145 GUIDE gs=0 - succeeded as much as you can in a closed dome.

Note 1.05 Gb of virtual memory used by IIF on lbtmu102.
top - 21:59:51 up 14 days,  2:01,  3 users,  load average: 4.29, 3.72, 3.13
Tasks:  84 total,   1 running,  83 sleeping,   0 stopped,   0 zombie
Cpu(s):  2.7% us,  1.0% sy,  0.0% ni, 96.2% id,  0.0% wa,  0.2% hi,  0.0% si
Mem:   1034616k total,   477104k used,   557512k free,    16064k buffers
Swap:  2621432k total,    20336k used,  2601096k free,   168940k cached

  PID USER      PR  NI  VIRT  RES  SHR S %CPU %MEM    TIME+  COMMAND
27889 telescop  16   0  308m  29m 6216 S  3.0  2.9   8:14.35 PCS
 5459 telescop  15   0  6292 2472 2232 S  2.7  0.2  69:25.63 gshmserver
28250 delapena  15   0 60404  34m  11m S  1.0  3.4   0:08.72 MCSGUI
 3538 root      15   0 20644 1584 1216 S  0.3  0.2   8:23.31 X
 5486 telescop  15   0  864m 191m 8932 S  0.3 19.0  33:22.61 IIF
    1 root      16   0  2220  464  436 S  0.0  0.0   0:00.86 init
    2 root      RT   0     0    0    0 S  0.0  0.0   0:00.27 migration/0
    3 root      34  19     0    0    0 S  0.0  0.0   0:00.03 ksoftirqd/0
    4 root      RT   0     0    0    0 S  0.0  0.0   0:00.35 migration/1
    5 root      34  19     0    0    0 S  0.0  0.0   0:00.04 ksoftirqd/1
    6 root       5 -10     0    0    0 S  0.0  0.0   0:00.06 events/0

22:06 pointto BS9152 ACTIVE gs=0

LBC tests

22:11 Change authorization to LBC

22:12 Turn on LBC

22:19 Play OB for Virgo Cluster

Apr  3 22:19:34 lbtmu102 LBT_PCS: Sample: 0 Time: 2454925.721921307966113 Theta: 2.701939894210525
Apr  3 22:19:35 lbtmu102 LBT_PCS: Sample: 0 Time: 2454925.721932881977409 Theta: 3.907799100806091
Apr  3 22:23:05 lbtmu102 LBT_PCS: Sample: 0 Time: 2454925.724363437388092 Theta: 3.892053786402228
Apr  3 22:23:19 lbtmu102 LBT_PCS: Sample: 0 Time: 2454925.724513900466263 Theta: 3.891054487827144
Apr  3 22:26:07 lbtmu102 LBT_PCS: Sample: 0 Time: 2454925.726458345074207 Theta: 3.877875821401204
Apr  3 22:26:08 lbtmu102 LBT_PCS: Sample: 0 Time: 2454925.726481493096799 Theta: 3.877002381732904
Apr  3 22:27:15 lbtmu102 LBT_PCS: Sample: 0 Time: 2454925.727256956044585 Theta: 3.871600913709056
Apr  3 22:28:44 lbtmu102 LBT_PCS: Sample: 0 Time: 2454925.728287048637867 Theta: 3.864300456967821
Apr  3 22:29:48 lbtmu102 LBT_PCS: Sample: 0 Time: 2454925.729027789551765 Theta: 3.858961035115608
Apr  3 22:32:24 lbtmu102 LBT_PCS: Sample: 0 Time: 2454925.730833345092833 Theta: 3.845625887270691
Apr  3 22:32:26 lbtmu102 LBT_PCS: Sample: 0 Time: 2454925.730844919104129 Theta: 3.846274553539367
Apr  3 22:33:44 lbtmu102 LBT_PCS: Sample: 0 Time: 2454925.731759270653129 Theta: 3.839347968937010
Apr  3 22:35:01 lbtmu102 LBT_PCS: Sample: 0 Time: 2454925.732650474645197 Theta: 3.832480348403657
Apr  3 22:36:18 lbtmu102 LBT_PCS: Sample: 0 Time: 2454925.733541678171605 Theta: 3.825496260093446
Apr  3 22:38:54 lbtmu102 LBT_PCS: Sample: 0 Time: 2454925.735347233712673 Theta: 3.810983337434567
Apr  3 22:38:55 lbtmu102 LBT_PCS: Sample: 0 Time: 2454925.735358807723969 Theta: 3.810567734403640
Apr  3 22:41:30 lbtmu102 LBT_PCS: Sample: 0 Time: 2454925.737152789253742 Theta: 3.795634417914865
Apr  3 22:41:31 lbtmu102 LBT_PCS: Sample: 0 Time: 2454925.737164363265038 Theta: 3.795536451871090
Apr  3 22:44:07 lbtmu102 LBT_PCS: Sample: 0 Time: 2454925.738969918806106 Theta: 3.780353058088884
Apr  3 22:46:47 lbtmu102 LBT_PCS: Sample: 0 Time: 2454925.740810196846724 Theta: 3.763995684473350

22:53 Play same OB again with telescope already in the same field
Apr  3 22:52:18 lbtmu102 LBT_PCS: Sample: 0 Time: 2454925.744641215074807 Theta: 3.728156275103824
Apr  3 22:52:19 lbtmu102 LBT_PCS: Sample: 0 Time: 2454925.744664363563061 Theta: 3.727932238327623
Apr  3 22:53:15 lbtmu102 LBT_PCS: Sample: 0 Time: 2454925.745312511455268 Theta: 3.721622194613916
Apr  3 22:54:55 lbtmu102 LBT_PCS: Sample: 0 Time: 2454925.746469919104129 Theta: 3.710175574838096
Apr  3 22:55:48 lbtmu102 LBT_PCS: Sample: 0 Time: 2454925.747083344962448 Theta: 3.704015391877917
Apr  3 22:58:24 lbtmu102 LBT_PCS: Sample: 0 Time: 2454925.748888900503516 Theta: 3.685504533892284
Apr  3 22:58:25 lbtmu102 LBT_PCS: Sample: 0 Time: 2454925.748900474514812 Theta: 3.684584194747113

23:36 Turn off LBC

IRTC

23:38 Back to IRTC - pointto BS9152 gs=0 TelM="ACTIVE" - succeeded in the closed dome sense. GCSL found a false star in the dome flat acquisition field.

Oil Temperature Testing

23:47 from hbssys
Oil Reservoir 0.7 degC
Input Cooling 2.7
Output Cooling -1.6
A-Input Water -22.4
A-Output Water 1.5
B-Input Water 1.8
B-Output Water: 2.1
Ta 2.9

EL T_In 6.6, 7.5, 6.6, 6.4
EL T_Out 5.9, 5.2, 4.7, 6.2

AZ T_In 5.4, 6.2, 6.1, 6.0
AZ T_Out 6.4, 6.6, 5.0, 6.6

Outside Ambient from weather station -4.5 (slowly cooling all day)
Dewpoint from weather station -4.3

23:53 from ECS-HBS
Temperature Setpoint -2.12
Coolant Return Temperature -2.19

------------------- Air ------ Steel
C-Ring Upper 1.39 1.15
C-Ring Middle 3.16 1.84
C-Ring Lower 4.17 2.51
C-Ring Ext Left 1.54 1.40
C-Ring Ext Right 1.10 1.40
Wind B Upper 0.78 1.25
Wind B Middle Left 0.90 1.21
Wind B Middle Right 1.08 1.42
Wind B Lower Right 0.88 1.87
Mirror Cell Low Right 1.06 1.13

23:58 SX Cell
Ref Temp 1.26
Glass 0.103

23:59 DX Cell
Ref Temp 0.94
Glass 0.447

Were unable to locate laser thermal probe to confirm these temperatures.

00:30 Reduced HBS setpoint offset in ECS from -6.0 to -9.0 in lbt.conf since oil seems to be 5 degC too warm compared to the average C-Ring. And I think the lower C-Ring is being heated by the oil. Stop and restart ECS to absorb value (probably not needed). However, the setpoint didn't change. This is because the -4.33 dewpoint is holding the setpoint up at -2.33. So -2.33 degC setpoint is delivering oil to the pads at +6 degC which means that the offset should be -8 degC. I've changed it back to -8.0 degC. Should we remove the dewpoint limit from this calculation? I've asked Chris to change the dewpoint limit from dewpoint+2 to dewpoint-2.

Telemetry r5 testing

01:06 Stopped and restarted PCS, PMCL, PMCR (and ECS just a while ago) to test the new version of Telemetry that Tony made last night.

01:07 pointto BS9177 gs=0 TelM="ACTIVE"

01:12 raise right primary mirror

on lbtmu103
top - 01:12:22 up 20 days,  3:50,  2 users,  load average: 5.84, 5.47, 4.70
Tasks:  81 total,   1 running,  80 sleeping,   0 stopped,   0 zombie
Cpu(s): 18.9% us,  3.8% sy,  0.0% ni, 76.7% id,  0.3% wa,  0.2% hi,  0.0% si
Mem:   1034552k total,   990428k used,    44124k free,   289512k buffers
Swap:  2621432k total,    24468k used,  2596964k free,   260852k cached

  PID USER      PR  NI  VIRT  RES  SHR S %CPU %MEM    TIME+  COMMAND
28949 telescop  15   0  407m  28m 6100 S 39.2  2.8   2:56.73 PMCL
27368 telescop  16   0  317m  38m 4876 S  3.0  3.8  69:12.46 OSS
23627 telescop  16   0  7124 1372 1184 S  2.3  0.1  78:36.16 gshmserver
 2565 root      16   0  2600  468  404 S  0.3  0.0   1:09.35 irqbalance
29022 telescop  16   0  2600  868  692 R  0.3  0.1   0:00.17 top
    1 root      16   0  2896  496  468 S  0.0  0.0   0:01.71 init
    2 root      RT   0     0    0    0 S  0.0  0.0   0:00.30 migration/0
    3 root      34  19     0    0    0 S  0.0  0.0   0:00.77 ksoftirqd/0
    4 root      RT   0     0    0    0 S  0.0  0.0   0:00.28 migration/1

on lbtmu105
top - 01:13:22 up 20 days,  3:51,  2 users,  load average: 3.61, 3.73, 3.92
Tasks:  83 total,   1 running,  82 sleeping,   0 stopped,   0 zombie
Cpu(s): 17.3% us,  4.2% sy,  0.0% ni, 78.4% id,  0.0% wa,  0.2% hi,  0.0% si
Mem:   1034552k total,   985832k used,    48720k free,   234916k buffers
Swap:  2621432k total,    24492k used,  2596940k free,   144612k cached

  PID USER      PR  NI  VIRT  RES  SHR S %CPU %MEM    TIME+  COMMAND
 1399 telescop  15   0  408m  28m 6116 S 40.6  2.8   3:30.71 PMCR
24736 telescop  15   0  7420 1304 1116 S  1.3  0.1  81:37.29 gshmserver
30742 telescop  16   0  229m  24m 3616 S  0.3  2.4   9:02.03 ENV
    1 root      16   0  1828  492  464 S  0.0  0.0   0:01.39 init
    2 root      RT   0     0    0    0 S  0.0  0.0   0:00.35 migration/0
    3 root      34  19     0    0    0 S  0.0  0.0   0:00.74 ksoftirqd/0
    4 root      RT   0     0    0    0 S  0.0  0.0   0:00.26 migration/1
    5 root      34  19     0    0    0 S  0.0  0.0   0:01.04 ksoftirqd/1
    6 root       5 -10     0    0    0 S  0.0  0.0   0:00.04 events/0

on lbtmu107
top - 01:16:44 up 14 days, 11:12,  4 users,  load average: 1.06, 3.54, 4.62
Tasks:  85 total,   1 running,  84 sleeping,   0 stopped,   0 zombie
Cpu(s):  0.3% us,  0.3% sy,  0.0% ni, 99.0% id,  0.3% wa,  0.0% hi,  0.0% si
Mem:   1034552k total,   339612k used,   694940k free,   109500k buffers
Swap:  2621432k total,    59760k used,  2561672k free,    95416k cached

  PID USER      PR  NI  VIRT  RES  SHR S %CPU %MEM    TIME+  COMMAND
25165 telescop  15   0  6484 1516 1332 S  1.0  0.1  65:47.11 gshmserver
13802 telescop  16   0  353m  42m 7996 S  0.3  4.2   0:28.95 GCSL
16491 telescop  16   0  2908  872  692 R  0.3  0.1   0:00.15 top
    1 root      16   0  3312  496  464 S  0.0  0.0   0:03.67 init
    2 root      RT   0     0    0    0 S  0.0  0.0   0:00.43 migration/0
    3 root      34  19     0    0    0 S  0.0  0.0   0:00.67 ksoftirqd/0
    4 root      RT   0     0    0    0 S  0.0  0.0   0:00.47 migration/1
    5 root      34  19     0    0    0 S  0.0  0.0   0:00.35 ksoftirqd/1
    6 root       5 -10     0    0    0 S  0.0  0.0   0:00.10 events/0
    7 root       5 -10     0    0    0 S  0.0  0.0   0:00.12 events/1

on lbtmu102
top - 01:17:35 up 14 days,  5:19,  2 users,  load average: 3.33, 3.02, 2.95
Tasks:  80 total,   1 running,  79 sleeping,   0 stopped,   0 zombie
Cpu(s):  1.2% us,  0.5% sy,  0.0% ni, 97.3% id,  0.8% wa,  0.2% hi,  0.0% si
Mem:   1034616k total,   435408k used,   599208k free,    34408k buffers
Swap:  2621432k total,    20336k used,  2601096k free,   134248k cached

  PID USER      PR  NI  VIRT  RES  SHR S %CPU %MEM    TIME+  COMMAND
32309 telescop  15   0  308m  29m 6232 S  2.7  2.9   0:20.23 PCS
 5459 telescop  15   0  6292 2472 2232 S  1.0  0.2  72:29.79 gshmserver
 5486 telescop  15   0  887m 195m 8976 S  0.3 19.3  34:34.93 IIF
32403 telescop  16   0  2480  864  688 R  0.3  0.1   0:00.01 top
    1 root      16   0  2220  464  436 S  0.0  0.0   0:00.86 init
    2 root      RT   0     0    0    0 S  0.0  0.0   0:00.27 migration/0
    3 root      34  19     0    0    0 S  0.0  0.0   0:00.03 ksoftirqd/0
    4 root      RT   0     0    0    0 S  0.0  0.0   0:00.35 migration/1
    5 root      34  19     0    0    0 S  0.0  0.0   0:00.04 ksoftirqd/1
    6 root       5 -10     0    0    0 S  0.0  0.0   0:00.06 events/0
    7 root       5 -10     0    0    0 S  0.0  0.0   0:00.06 events/1

02:00 John finishes editting ptspiral.csh

02:10 Aaron testing the moving of swing arms with OSS/OSSGUI

02:12 OSSUI crashes in the middle of the process - needs more work. MCSPU complains that Dynamic Balance has stow pins disabled. We think this is because the retract of Right PF swing arm started before the crash. Also note that the StowPin position swtiches are not at 90 deg but at 90.11 deg based on tonight's encoder reset at horizon.

02:30 Wind still gusting to 25 m/sec with 100% humidty. At -6 degC, this means blowing snow.

-- JohnHill - 02 Apr 2009
Topic revision: r5 - 04 Apr 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback